OpenStudy (anonymous):

Describe a real world scenario that is an arithmetic series. Show your work as you find the 7th term of your scenario.

OpenStudy (anonymous):

help

OpenStudy (jim766):

George had $10 in a bank account. Each week he added $5. How much would be in the account in 7 weeks.

OpenStudy (jim766):

An = a1+ (n-1)d A7 = 10 + (7-1)5